### v1

# Language-Driven Closed-Loop Grasping with Model-Predictive Trajectory Replanning

Abstract
Combining a vision module inside a closed-loop control system for a seamless movement of a robot in a manipulation task is challenging due to the inconsistent update rates between utilized modules. This task is even more difficult in a dynamic environment, e.g., objects are moving. This paper presents a modular zero-shot framework for language-driven manipulation of (dynamic) objects through a closed-loop control system with real-time trajectory replanning and an online 6D object pose localization. We segment an object within by leveraging a vision language model via language commands. Then, guided by natural language commands, a closed-loop system, including a unified pose estimation and tracking and online trajectory planning, is utilized to continuously track this object and compute the optimal trajectory in real-time. Our proposed zero-shot framework provides a smooth trajectory that avoids jerky movements and ensures the robot can grasp a non-stationary object. Experiment results exhibit the real-time capability of the proposed zero-shot modular framework for the trajectory optimization module to accurately and efficiently grasp moving objects, i.e., up to update rates for the online 6D pose localization module and update rates for the receding-horizon trajectory optimization. These advantages highlight the modular framework’s potential applications in robotics and human-robot interaction; see the video in https://www.acin.tuwien.ac.at/en/6e64/.
I INTRODUCTION
Integrating artificial intelligence (AI) in robotic systems transforms human-robot collaboration, particularly in dynamic and complex environments. Traditional robotic systems typically rely on pre-programmed tasks and controlled environments with limited variability. However, real-world applications demand higher adaptability and precision, especially when robots need to interact with humans in real-time. Imagine we want an assistant robot to grasp a hammer among a clutter of workshop objects such as scissors, scales, screwdrivers, or handrails. The robot must be capable of continuously detecting and grasping the hammer, even if the hammer can be moved. While humans can naturally understand how to grasp a hammer when instructed, it is still challenging for robots to identify precise grasping actions for objects based on natural language commands, a task known as language-driven grasp detection [1]. This simple example encompasses three challenges, i.e., natural language understanding, real-time object localization, and real-time trajectory planning. Despite recent advancements, bridging the gap between language, vision, and control in real-world robotic experiments remains a challenging task [2]. Furthermore, recent works, e.g., LLaRP [3], Open X-embodiment [4], often focus on an end-to-end approach relying on pre-defined large skill datasets. This reliance on individual skill acquisition is usually seen as a significant bottleneck for application in various systems, primarily due to mismatches in dynamic models, which result in the jerky motion of the robot. Furthermore, these end-to-end approaches often neglect kinematic and dynamic constraints, leading to unsafe and inaccurate robot motions. These requirements are significant concerns for meeting industrial standards, where precise and reliable operations are critical. Thus, in this work, our focus is creating a zero-shot modular framework utilizing a vast knowledge of foundation models for vision-language understanding and smoothly controlling the robot at a granular action level, bypassing the need for extensive data collection or manual programming of individual basic movements. Also, with the modular fashion, we can integrate these individual system constraints into the trajectory optimization process to ensure that the robot’s movements are safe and precise.
To address the challenge above in language-driven grasping tasks, we propose a modular zero-shot framework comprising (i) a vision-language model to analyze the human prompts and to ground the mentioned object, (ii) a vision module to localize the object in real-time, and (iii) the receding-horizon trajectory optimization to execute a smooth trajectory toward the object. Specifically, we first utilize the open-vocabulary detectors such as NanoOWL [5] and OWL-ViT2 [6] leveraging contrastively pre-train image and text encoders on large-scale image-text data RegionCLIP [7]. This allows us to extract spatial-geometric information of relevant objects from natural language commands. Next, we employ the unified foundation model for 6D object pose localization, FoundationPose [8], to efficiently estimate and continuously track the object’s 6D pose. Given that the object is not stationary, we employ our recent work, the model-predictive trajectory optimization (MP-TrajOpt)[9], to dynamically guide the robot toward the object. An overview of the proposed framework is illustrated in Figure 1, which is presented in detail in Section III.
The remainder of the paper is as follows. After discussing related work in the next section, Section III presents the proposed framework, including three modular modules. Subsequently, experiments and analysis, including comparisons with other approaches and limitations, are discussed in Section IV. Finally, Section V concludes this work and gives some remarks.
II Related Work
Foundation Models for Robotic Manipulation. Various methods have been proposed to integrate foundation models into robotic manipulation [3, 10]. For instance, Ha et al.[11] use large language models (LLMs) as knowledge bases to decide tasks and utilize sampling-based robot planners to create comprehensive manipulation skill datasets. These datasets can then be used to train diffusion policies for diverse manipulation tasks using natural language commands. While the trend of applying foundation models to robotics is growing [12], it still faces significant uncertainties [13]. One major issue is the need for improved functional competence in these models [13], which limits their ability to tackle novel planning problems that require an understanding of real-world dynamics [14]. Additionally, enabling language models to perceive physical environments often involves providing textual descriptions and incorporating vision during decoding [15]. To close the perception-action loop, an embodied language model must also be capable of action, typically achieved through a library of predefined primitives. However, these primitives vary across robotic platforms, often requiring a large dataset to generalize effectively. Unlike previous methods, we leverage pre-trained language models for their extensive open-world knowledge and focus on the complex task of planning for dynamic object manipulation. Our approach incorporates system dynamic constraints, enabling a seamless integration of perception and action. This method addresses the limitations of previous models by enhancing functional competence and facilitating real-time interaction and adaptation.
Language-driven object detection. Traditionally, language-driven object detection methods were designed for limited categories, a.k.a. closed-vocabulary object detection, like DETR [16]. Recently, using large vision-language models (VLMs) like CLIP [17] has opened doors to open-vocabulary object detection. Unlike traditional methods, open-vocabulary object detection methods OWLv2 [6] and RegionCLIP [7] are not restricted to predefined categories. This allows them to excel in tasks like zero-shot object detection and segmentation based on natural language descriptions. This shift towards open-vocabulary object detection [6] extends beyond just identifying objects. It has applications in related areas like affordance detection [18], where the focus is identifying potential actions that can be performed on objects. Inspired by OWLv2, we utilize this method in our framework to enable zero-shot object detection via natural language commands.
Object pose localization. Recent advancements in object pose estimation have focused on developing techniques that handle novel objects without requiring retraining, addressing a key challenge in applying deep learning models to industrial settings. Nguyen et al. [19] introduced GigaPose, simplifying the pose estimation process to a single correspondence match between a query image and a set of templates. More recently, Wen et al. [8] presented FoundationPose, a unified 6D object pose estimation and tracking model, utilizing large-scale synthetic training data and a novel transformer-based architecture. FoundationPose supports both model-based and model-free setups with novel objects. However, manual segmentation of objects in the first frame from pre-recorded videos is still required. Therefore, in this work, we integrate a language-driven object detection module that could enable end-to-end applications from human commands to robotic grasping actions in a closed-loop system.
Real-time trajectory optimization. In recent years, significant interest has been in extending trajectory optimization to online planning using a receding horizon. This interest spans both gradient-based methods [20] and sampling-based approaches [21]. The sequence-of-constraints Model Predictive Control (MPC) introduced by Toussaint et al. [22] addresses task-and-motion planning (TAMP) in three phases: initially, task planning generates waypoints; subsequently, the timing of these waypoints is optimized to form a reference trajectory; finally, MPC tracks this reference trajectory to compute collision-free paths over a short planning horizon. Note that this approach requires a global reference to incorporate waypoints into the MPC. Tackling this issue, in our recent work [9], we propose the model-predictive trajectory optimization (MP-TrajOpt) that avoids calculating a reference trajectory for waypoint timing. Instead, this MPC approach employs a cost-to-go analysis for waypoints. The advantages of integrating MP-TrajOpt are twofold. First, to complete a general grasping task, the robot must navigate specific waypoints to avoid collisions between its gripper and the object. For example, in a top-down grasp, the robot should pass through a waypoint located above the object along the -axis, as illustrated in Figure 3. Second, passing waypoints with non-zero velocity allows for a more natural and smooth grasping action, which differs from most state-of-the-art methods [23, 24].
Leveraging vast knowledge from foundation models, we tackle the language-driven task in this world. Our contributions are summarized as follows:
We present a zero-shot modular framework for end-to-end language-driven closed-loop grasping tasks that operate without training data. It is important to note that the primary goal of this work is to offer a comprehensive implementation rather than serve as an evaluation benchmark.
We provide extensive demonstrations with real-world objects and comparisons with state-of-the-art methods [24]. Additionally, we open-source our implementation at language-driven-closed-loop-grasping.github.io to support community development.
III Methods
In this section, we explain the three main modules: language-driven object detection, real-time object pose localization, and online trajectory optimization (MP-TrajOpt), illustrated in Figure 1. Consider a scenario in a workshop where a human instructs a robot to grasp a tool with a known CAD model and place it in a predefined location. First, the language-driven object detection module processes the image to determine the 2D location of the tool and generate a binary mask. This mask, combined with the CAD model, is used by the real-time object pose localization module to estimate the initial pose of the tool. The pose is then refined using a linear Kalman filter for smoothness. Finally, the online trajectory optimization module plans and executes the grasp and placement actions, ensuring smooth and collision-free movements.
III-A Language-Driven Object Detection
To determine the 2D location of a target object specified in a natural language command, such as ”Grasp the orange drill,” we first extract key phrases from the command, in this case, ”orange drill.” Concurrently, an RGB image is input into the OWLv2 model [6] to generate pseudo-box annotations for image-text pairs, as illustrated in Figure 2. Although we utilize the pre-trained weights of the OWLv2 model without additional training, this approach is both efficient and effective at pinpointing the object’s center within the image. In comparison to more complex segmentation methods, such as SAM [25] or CNOS [26], which require several seconds for inference, the OWLv2 model can ground the object within across all experimental cases. Upon obtaining the bounding box of the target object, a binary mask is generated. This binary mask serves as the input for the subsequent object pose localization process.
III-B Object pose localization module
Inspired by the state-of-the-art pose estimation method by Wen et al.[8], we present a real-time object pose estimation and tracking framework employing a three-stage pipeline, as depicted in the schematic of Figure 2.
In the pose initialization stage, we align and synchronize data from the depth sensor and the RGB sensor of an RGB-D camera. Subsequently, leveraging the object detection mask generated by the method mentioned earlier, we formulate an initial coarse pose estimation. This initial estimation serves as input to a pose refinement network in the subsequent stage. The architecture of this refinement network, depicted in the second stage of Figure 2, is designed to infer translation updates and rotation updates of the corresponding object. Each update is individually processed through a transformer encoder[27] and then linearly projected to the output dimension. Notably, both and are expressed relative to the camera frame. The initial coarse pose is refined utilizing the following update equations [8]:
where indicates the update operation on . This disentangled representation approach effectively decouples the translation update from the updated orientation. The refinement network is trained under supervision [8] using the L2 loss function:
where and denote the ground truth values, and and are empirically determined weights used to balance the loss terms, both set to .
Following the refinement of pose hypotheses, a hierarchical pose ranking network computes their respective scores in the pose selection stage of Figure 2. Employing a hierarchical comparison and contrast validation approach proposed in[8], the network computes the loss for each pair of hypotheses, ultimately selecting the final pose estimate with the highest score. Notably, this network employs the same backbone architecture for feature extraction as the refinement network but with slight modifications by adding the multi-head self-attention. The extracted features are concatenated, tokenized, and processed through a multi-head self-attention mechanism and a pose ranking encoder to generate a feature embedding across all pose hypotheses. Then, multi-head self-attention is applied to the concatenated feature embedding , capturing pose alignment information from all hypotheses. Finally, the attended features are linearly projected to produce the scores , which are associated with the pose hypotheses. By selecting the highest scores for the pose hypotheses, we obtain accurate and meaningful pose estimations, ensuring the reliability of the final pose output.
To enhance the accuracy and smoothness of pose estimation ] of the object, we incorporate a simple Kalman Filter (KF) [28] of the state vector
with is the vector of three Euler angles computed from the object’s orientation [29]. We consider the process model for the LKF with the update time and at the time index in the following
with and is white Gaussian noise. Note that is the Kronecker product, and is the identity matrix with the size . The measurement model is simply
with and is normal distributed measurement noise. We consider the following steps
to update while eliminating jerky measurements from the pose selection stage. Note that and are the user-defined process and measurement noise covariance, respectively.
III-C Model-predictive trajectory optimization (MP-TrajOpt)
III-C1 Mathematical model
The rigid-body dynamical model of a fully actuated robot is given by
with the generalized coordinates , the positive definite mass matrix , the matrix of Coriolis and centrifugal terms , the gravitational forces , and the generalized torques . We consider only a rigid-body model for simplicity of presentation, neglecting elastic joints. For the full elastic joint model, the reader is referred to [30] for a detailed derivation of the model and the controller.
The Cartesian end-effector pose of the robot, obtained by the forward kinematics of the robot, is described as a homogeneous transformation and given by
containing the position and the rotation matrix and denotes a zero matrix of matching size.
The RH-TrajOpt from [9] requires a goal and waypoints in the joint space. We obtain those joint space solutions from the object pose localization using an analytic inverse kinematics solution of the KUKA LBR iiwa 14 R820 [31]. Compared to numerical methods, the computational efficiency of the analytical solution is much better. Fast computation time is crucial for real-time applications.
III-C2 Control
We compensate the nonlinear robot dynamics with the computed torque state-feedback law
which simplifies the dynamic model (8) to the double integrator system
with the new control input . The remaining linear dynamics (11) are stabilized using
where and are suitable gain matrices which render the closed-loop trajectory error dynamics exponentially stable. The desired reference trajectory in the joint space is given by with its time derivatives and and the time .
III-C3 Planning
The controller (12) requires at least two times continuously differentiable reference trajectory from the planner. We apply triple integrator dynamics for the planning algorithm for additional smoothness. The state vector is defined as with the input . Assuming piecewise-linear inputs with the sampling time leads to the first-order-hold discrete-time state-space formulation
where
The symbol denotes the Kronecker product, and is the identity matrix of size .
We summarize the model-predictive trajectory planning (MP-TrajOpt) approach from our recent work [9]. We calculate the inverse kinematics solution [32] of the pre-grasp and grasp pose and the pre-place and place pose, respectively, to obtain a joint space waypoint and goal configuration , and . To consider the waypoint along the way to the goal, we split the planning horizon into two parts, namely and as soon as the waypoint is reachable within the horizon. If the waypoint is reachable, the planning algorithm introduces a constraint for the waypoint to pass it with a certain tolerance. Similarly, the planner introduces an endpoint constraint for the goal when it is reachable within the horizon.
The optimal system state and input trajectory for the -th iteration of the planner are obtained from the discrete-time optimization problem
s.t. | ||||
the cost terms and drive the trajectory to the waypoint and the goal, respectively. The regularization cost on ensures smooth trajectories, and is a collision avoidance term. The weights are tuning parameters. The trajectory must adhere to the system dynamics (15b) with the initial states obtained by the previous planner iteration (15c). The endpoint is constraint to an equilibrium point in (15d), and the states and inputs are bounded using upper , and lower , bounds. The endpoints of the split horizon are constrained to a waypoint tolerance set and a goal tolerance set , respectively. For details on computing the horizon lengths and , refer to [9].
We use a state machine to guide the robot through the pick-and-place task. First, the robot must approach the object to grasp it. We derive a grasp pose and a pre-grasp waypoint from the object pose. The pre-grasp pose is placed above the object to ensure a proper approach direction and a collision-free grasp. Figure 3 illustrates a simple example of this pre-grasp waypoint. Note that moving the object further than a tolerance dynamically changes these poses. When the robot end-effector reaches the grasp pose, the robot grasps the object and can plan toward the place location. Similarly to the grasp poses, a pre-place and place pose are defined to place the object.
IV Experiments
Experiments are carried out on grasping tasks in real-world scenarios with a 7-DoF KUKA LBR iiwa 14 R820 and novel objects, captured in Figure 4. We utilize the D435i RealSense camera in an eye-to-hand setup where this camera observes the robot within its workspace like in Figure 3.
Three modules of the proposed framework, see Figure 1, are processed on a desktop with Intel Core i7 10700K and 32 GB of RAM. Inference processes of the language-driven module and the object pose localization are run on the NVIDIA RTX 3080 with 8GB of VRAM. The optimization problem described in Equation (15) is implemented as a Python-based ROS node utilizing CasADi [33]. It is solved using the nonlinear interior point solver IPOPT [34] and the MA57 linear solver, all running on an Ubuntu 22.04 system. This implementation achieves planning times under , which includes the online resolution of the analytic inverse kinematics for new Cartesian waypoints and a desired goal [31]. To ensure compatibility of inverse kinematics solutions for these waypoints, the solution nearest to the previous one is selected using a least-squares criterion.
IV-A Statistical results of vision modules
Considering the same test setup depicted in Figure 3 and all ten objects shown in Figure 4, we conducted statistical analysis on the inference speeds and success rates of using another state-of-the-art language-driven vision module, namely CNOS [26] with FastSAM [25], in conjunction with the presented object localization module. For all test cases, the prompt ”Please grasp the [object]” was used, with the placeholder [object] being replaced by terms such as ”orange drill,” ”blue block,” ”orange block,” ”green block,” etc. Each object underwent five trials. Table II presents the computation times of the inference processes of the vision modules using different GPUs. The term ”oFP & LKF” refers to the online version of FoundationPose with a linear Kalman filter, which is utilized in our object pose localization module. It is noteworthy that the CNOS & FastSAM models are pretty slow, requiring several seconds to segment objects through language commands. Conversely, the lightweight OWLv2 requires approximately to compute a binary mask of the segmented object via a prompt. It is not surprising that with a Nvidia RTX 4090, an update rate of approximately can be achieved using the language-driven module employing OWLv2 [6], and approximately with the proposed object localization module. Under identical lighting conditions and prompts across 50 trials, OWLv2 detected objects in 46 instances, whereas the other methods succeeded in only 40 cases. It is important to note that once an object is detected and the binary mask is provided, the object localization model can reliably track the object.
IV-B Qualitative results of the performance of the object localization module
The qualitative performance of the object localization module was evaluated using ground truth data from an OptiTrack system. This assessment aimed to verify the accuracy of the object localization module. Figure 5 presents the tracking results for the blue Lego block (depicted in Figure 4). The 3D position trajectory of this Lego block closely matches the measurements obtained from the OptiTrack system. It is important to note that these measurements are transformed to the camera origin, approximately away from the object. The maximum positional error observed is approximately , representing 2% relative to the distance to the camera. The orientation error is constrained, with a maximum error of in the roll angle. This evaluation is influenced by the jerky movements of the human hand while holding the object and the intense IR light emitted by the OptiTrack system. The errors are significantly lower when the object remains stationary, as illustrated in Figure 5. Notably, the vision module demonstrates promising accuracy in dynamic scenarios and can meet industrial standards in static scenarios.
IV-C Robotic experiments
We evaluated ten novel objects, depicted in Figure 4, using our proposed modular framework, as well as AnyGrasp [35], LGD [36], and VoxPoser [24]. Each object was subjected to 5 trials. It is important to note that only our framework supports catching moving objects; the other three methods were tested in static scenarios.
AnyGrasp [35], which does not utilize natural language prompts, achieved 37 successful trials out of 50 without prompts. The LGD [36] approach encountered 27 failures due to incorrect object detections in cluttered scenarios. Unlike LGD [36] and AnyGrasp [35], both VoxPoser [24] and our modular framework consider the pre-grasp pose for all cases, as shown in Figure 3.
VoxPoser utilizes the standard task space controller [37], which directs the robot toward the segmented object in the scenario. However, due to segmentation errors and occasional singularities in the task space controller, VoxPoser only achieved 33 successful cases. In contrast, our approach encountered only eight failures, primarily due to segmentation errors from OWLv2. Additionally, our proposed framework demonstrated the capability to replan in all 50 trials within dynamic scenarios.
Figure 6 shows overlay images from two experiments with our proposed framework. Small windows on the upper right at each image indicate the view from the Realsense camera. In Figure 6 (a), with the given text prompt: “Grasp the plier” and although the plier is impeded in the scenario, the proposed framework still outputs the correct pose of the plier, see the small window on the upper right of Figure 6 (a). The proposed framework generally exhibits good results in localizing, tracking, and accurately grasping the object in a closed loop. More demonstrations are shown on the project page language-driven-closed-loop-grasping.github.io.
Hardware | Success rate |
AnyGrasp [35] | 0.74 | ✗ | ||
LGD [36] | 0.46 | ✗ | ||
VoxPoser [24] | 0.66 | ✗ | ||
Ours | 0.92 | ✓ |
IV-D Limitations
Although we achieve good results, typical failure cases arise from a combination of challenges, including texture-less surfaces, severe occlusion, and limited edge cues, which lead to incorrect object detection and subsequent failed object localization. Additionally, the inference speed of the object localization process becomes a bottleneck when using low-profile GPUs, such as the NVIDIA RTX 3080. While upgrading to a more powerful GPU is an option, it is not always feasible, particularly in low-budget setups. Alternative solutions include incorporating prediction modules, such as those proposed by [38], or developing compact models for the object localization module. Furthermore, the proposed zero-shot framework cannot re-identify the object when object localization fails due to occlusions or rapid movement. Therefore, an upgraded framework with re-identification capabilities must address these issues.
V Conclusions
We have introduced a novel zero-shot modular framework designed to address the challenges of language-driven manipulation of dynamic objects within a closed-loop control system. Our framework leverages a vision-language model for natural language processing and real-time object segmentation, allowing for the seamless integration of 6D object pose localization and receding-horizon trajectory optimization to achieve precise and smooth robotic movements. The intensive experimental results highlight the robustness and efficiency of our framework, achieving a 92% success rate and superior inference speeds compared to state-of-the-art methods. Furthermore, the object localization module demonstrated high reliability with minimal positional and orientation errors, particularly in dynamic environments requiring real-time replanning. Future work will focus on further improving 6D pose estimation by continuously localizing object segmentation to prevent tracking loss and detecting affordance grasping objects through a language-driven approach.